# Benoit Decerf [^1] World Bank    Guillaume Haeringer [^2] Baruch College    Martin Van der Linden [^3] Independent scholar

### 07 Jan 2024 v1

@@ Abstract
Mandatory clearinghouse participation is a standard setting in many school districts, and students can appeal (block) their assignments only with a valid reason. When students do not observe the assignments of other students, claims relying on justified envy are hard to establish. A student’s appeal is said to be legitimate if there is no profile of preferences of the other students that can rationalize the student’s assignment, and an assignment is incontestable if no student has a legitimate complaint. Incontestability is shown to be equivalent to individual rationality, non-wastefulness, and respect for top-priority sets. This latter property is a weakening of non-justified envy that relates to Hall’s marriage theorem. We show that any stable mechanism is incontestable, as well as any mechanism that Pareto dominates a stable mechanism. The Top-Trading Cycle mechanism and some of its variants are incontestable (but Boston is not). Incontestable mechanisms are $i$-indinstiguishable (Li,, 2017), and share many similar incentive properties.
@@ 1 Introduction
Stability is a foundational concept in matching theory, serving as one of the primary objectives for both matching problems (e.g., doctors and hospitals) and assignment problems (e.g, school choice). While stability is unquestionably crucial for matching problems, we contend that its justification is less clear for assignment problems like school choice. One rationale for stability is its normative appeal, as it encapsulates a notion of fairness. However, while undoubtedly appealing, fairness comes at a cost because in assignment problems stability does not guarantee efficiency (Ergin,, 2002). Another rationale for stability is that it helps preventing market failures: unstable matchings are ‘blocked,’ which may in turn undermine market participation —see Roth, (1991) and Kagel and Roth, (2000). [^4] It should be noted, though, that stability does not always prevent market failure. See for instance McKinney et al., (2005)., [^5] The rationale is easy to understand when considering a centralized job market, where workers can offer themselves to firms that rank higher in their preferences after being notified of their matches. If the matching were not stable, some firms might accept these ‘decentralized’ offers and thus reject candidates they have been assigned to by the clearinghouse. Incentives to participate in the matching mechanism are then weakened, which may end up in market failure. Note that this intuition aligns well with scenarios where market participation is voluntary. However, in settings where assignments can only be decided by a central authority and/or a clearinghouse (like in most school districts), blocking pairs may not be allowed to alter assignments. Why stability matters in such cases seems thus less immediately evident. [^6] Note that under the current National Resident Matching Program the matching produced by the clearinghouse is also binding —see Curtin and Signer, (2019).
This paper’s objective is to develop a notion of stability more tailored to school choice contexts by revisiting the concept of ‘blocking.’ In such markets, blocking takes the form of an appeal submitted to the school district. However, in most school districts students can only successfully appeal if they provide a valid reason, that is, a justification for why their school assignment should be different than the one they have been given. A school choice assignment can thus be deemed ‘stable’ if it is ‘appeal-proof.’ We assert that what can constitute an appeal-proof assignment depends on what information is available to the students. To understand our point, consider the information requirements for the three conditions that define together stability: individual rationality, non-wastefulness, and absence of justified envy. Individual rationality requires no information beyond a student’s own assignment. Non-wastefulness demands knowledge of school enrollment sizes (usually publicly available information) in order to formulate a legitimate appeal. The absence of justified envy, however, implicitly implies that students know the assignments of the other students. This is problematic because students are usually unlikely to have that information, at least in the first days or weeks after they are notified of their assignments (which we implicitly assume in this paper). Without information on who is assigned to a particular school, complaints of priority violations are challenging to establish.
Our contribution starts by formalizing the information held by students when participating in a school choice mechanism in a similar fashion as Fernandez, (2018) or Chen and Möller, (2023). Unlike those two papers, we consider both the interim and ex-post stages. [^7] Fernandez, (2018) or Chen and Möller, (2023) only consider ex-post information structures. In the interim case, each student only knows the set of schools (and their priority rankings and capacities), the set of students, and her own preferences (i.e., the preference list submitted to the mechanism). In the ex-post case, students also know their assignment (but not the assignments of the other students), and how many students have been assigned to each school. Because we aim at capturing a notion of blocking when information is not complete, our starting point is to consider the ex-post stage. We say that a student has a legitimate complaint about her assignment if the information she has is enough to show that her assignment cannot be the result of a stable assignment, that is, if there is no preference profile of the other students such that her assignment is part of a stable assignment for that profile. (So, stable assignments are trivially appeal-proof.) Our approach to legitimate complaints is thus akin to a ‘worst case scenario,’ similar in spirit to the blocking concept developed by Liu et al., (2014) for matching problems with incomplete information. Our notion is also deliberately weak, in the sense that a student’s complaint in our setting does not require a student to identify the school she should be assigned to. However, our characterization result will entail that a legitimate complaint necessarily implies that the student must be assigned to a more preferred school. An assignment where no student has a legitimate complaint is called incontestable.
@@
Given our definition, checking whether an assignment is incontestable may thus require to consider, for each student, all possible stable assignments given all possible profile of preferences of the other students. We show that incontestability is in fact much easier to verify: an assignment is incontestable if, and only if it is individually rational, non-wasteful and satisfies a novel property that we call respect for top-priority sets (Theorem 1). That property, which can be understood as a non-justified envy condition when students cannot observe or deduce the other students’ assignments, only applies to students who have a set of schools that they prefer to any other school and that cannot be filled only with students who have higher priority. We call such a set a top-priority set, and an assignment respects top-priority sets if any student with such a set is assigned to a school in that set. We show that when that property is violated for a student, one of the schools in her top-priority set must either be under capacity or have admitted a student with a lower priority. [^8] Respect for top-priority set is related to the failure of Hall,’s (1935) marriage condition. See Section 3.2 for more details.
@@
Theorem 1 implies that a student will not have a legitimate complaint whenever she is assigned to a school in her top-priority set (if she has such a set), or whenever she is assigned to any individually rational outcome (if she does not have such a set), and all schools preferred to her assignment are full. Under a mild consistency property, called top-top consistency, we obtain a characterization of incontestable mechanisms’ outcomes that parallels the characterization offered by Theorem 1. Top-top consistency is a weakening of the standard consistency property that applies only for student-school pairs that are mutually ranked first.999 We show in the Appendix that the most common mechanisms considered in the literature are top-top consistent. The characterization we obtain is the following (Theorem 2). For students with a top-priority set, any school in that set can be the outcome of an incontestable and top-top consistent mechanism. That is, for any school in that set, and for any incontestable and top-top consistent mechanism, there exists a preference profile of the other students under which the student is assigned to that school. As for the students without a top-priority set, the set of possible outcomes is the set of all individually rational assignments. Remarkably, this characterization does not depend on the particular mechanism that is used. In other words, mechanisms that are incontestable and top-top consistent are $i$-indistinguishable in the sense of Li, (2017).
Another remarkable property of incontestability is its relation to Pareto improvements: any assignment that Pareto dominates an incontestable assignment is also incontestable. Since stable assignments are always incontestable, any mechanism that Pareto improve the student-optimal stable mechanism is also incontestable. This implies first that efficiency and incontestability are not incompatible, and that any mechanism designed to Pareto improve the Student-Optimal Stable mechanism (SOSM) is incontestable. One such mechanism is Kesten,’s (2010) Efficiency Adjusted Deferred Acceptance Mechanism (EADAM). There are also incontestable mechanisms that are not Pareto comparable to SOSM. The Top-Trading Cycle mechanism (and some of its variant) is also incontestable (but the Boston mechanism is not). It is worth noting that the weakenings of the stability concept proposed in the literature do not include TTC outcomes as possibly ‘stable.’
@@
Our characterization in Theorem 2 turns out to be extremely useful to study (and compare) the incentive properties of incontestable mechanisms. Of course, such mechanisms are not guaranteed to be strategyproof. For instance, EADAM is not strategyproof. Reny, (2022) nevertheless shows that under EADAM, truth-telling is a maxmin optimal strategy. Theorem 2 makes it relatively easy to extend Reny’s theorem to any incontestable and top-top consistent mechanism. We also show that we can easily get results that compare the manipulability of mechanisms, when varying the maximal number of schools students are permitted to include in their submitted preference lists.
@@
Related literature   Our paper intersects with several branches of the matching and assignment literature. The first of these branches deals with proposals to obtain assignments that Pareto dominate the student-optimal stable mechanism. In broad strokes, those mechanisms allow for certain students to have justified envy. [^10] See Dur et al., (2019) who propose a family of mechanisms based on this principle. A key contribution in that direction is Kesten,’s (2010)’s EADAM algorithm. His mechanism consist of identifying ‘interrupters,’ that is, students who will make the student-optimal assignment inefficient because they prefer certain schools to their assignment. In Kesten’s algorithm, students are first asked whether they want to waive their priorities at the schools for which they are interrupters. Ehlers and Morrill, (2020) and Reny, (2022) propose different rationales than Kesten’s, but end up with a mechanism that is outcome equivalent to EADAM (when all students waive their priorities). Our approach differs from these contributions in that our notion of blocking is entirely based on the information available to the students. It is the lack of complete information on students’ side that allows us to implement assignments that Pareto dominates the student-optimal assignment, and not a tolerance for some level of justified envy.
@@
We are not the first ones to propose a definition of stability that depends on the information available to participants. An early contribution in that respect is Chakraborty et al., (2010) who consider the problem where agents can infer the types of potential partners from the observation of the whole matching. Another contribution is Yenmez, (2013), who considers ex-ante and interim stability conditions. [^11] See also Ehlers and Massó, (2015). A notable contribution is Liu et al., (2014), who extend Crawford and Knoer,’s (1981) model where firms do no observe workers’ types. Our work is similar to this latter in the sense that blocking and legitimate complaint are both based on a worst-case scenario principle. In Liu et al., (2014), a worker-firm pair blocks a matching if they can be both better off for any possible belief, and in our paper a legitimate complaint is a complaint that holds for any possible preference profile of the other students. But our contribution is somehow orthogonal to Liu et al., (2014) or Chakraborty et al., (2010). They assume that firms observe the whole matching but not workers’ types, whereas students in our model observe schools’ priorities (i.e., their ‘types’) but not the complete assignment.
@@
Our paper also deals with the question of agent’s information when participating in a mechanism, and Li,’s (2017) contribution is obviously key to frame and interpret some of our results. Our results complement those of Möller, (2022), who analyze conditions under which students can be guaranteed that the central authority has used the mechanism it initially announced. Möller,’s (2022) results and ours share some similarities but are obtained under different approaches. Möller,’s contribution is motivated by questioning the central authority’s credibility, while our initial motivation is to understand what stability entails in a low information environment. [^12] Möller, assumes that all students observe the complete assignments. See also for Hakimov and Raghavan, (2022) for another contribution in a similar vein.
@@ 3.1 Information and complaints
It will prove convenient to make the distinction between the information held at interim and ex-post stages of an assignment mechanism, that is, before or after the assignment mechanism has computed an assignment for all students. [^13] We will not consider in this paper the ex-ante stage (i.e., before students know their preferences over schools). We thus say that an information set $\mathcal{H}_{i}$ of a student $i$ is interim if $\succ_{i}\in\mathcal{H}_{i}$ and $\mathcal{H}_{i}$ does not contain any (partial) information about an assignment. The information set $\mathcal{H}_{i}$ is ex-post if $\succ_{i}\in\mathcal{H}_{i}$ and $\mathcal{H}_{i}$ contains some information about a realized assignment.
@@
A school choice problem $\Gamma=(I,S,\succ,r,q)$ is compatible with an interim information set $\mathcal{H}_{i}$ if there is no element in $\Gamma$ that contradicts an element in $\mathcal{H}_{i}$. A school choice problem $\Gamma=(I,S,\succ,r,q)$ and assignment $\mu$ are compatible with an ex-post information set $\mathcal{H}_{i}$ if there is no element in $\Gamma$ and $\mu$ that contradicts an element in $\mathcal{H}_{i}$.141414 We restrict in this paper to situations where each student is not informed about the other students’ preferences and assignments. So $\Gamma=(I,S,\succ^{\prime},r,q)$ is compatible with $\mathcal{H}_{i}=\{I,S,\succ_{i},r,q\}$ if $\succ_{i}^{\prime}=\succ_{i}$. If $\mu(i)\in\mathcal{H}_{i}$ then $\mu^{\prime}$ is compatible if $\mu^{\prime}(i)=\mu(i)$.
@@
We say that a student $i$ with an ex-post information set $\mathcal{H}_{i}$ has a legitimate complaint at an assignment $\mu$ if $\mathcal{H}_{i}$ is enough to show that if $\mu$ were a stable assignment then $i$ should not be assigned to $\mu(i)$. To understand more formally what we mean by the fact that $\mathcal{H}_{i}$ is ‘enough,’ consider its negation. If the information set $\mathcal{H}_{i}$ is ‘not enough’ for $i$’s complaint then it must be that whether $i$ can be assigned to $\mu(i)$ at a stable assignment depends on some elements not included in $\mathcal{H}_{i}$. That is, there are two problems, say, $\widehat{\Gamma}$ and $\widetilde{\Gamma}$, that are both compatible with $\mathcal{H}_{i}$, and there exists an assignment $\widehat{\mu}$ that is stable for $\widehat{\Gamma}$ such that $\widehat{\mu}(i)=\mu(i)$ but there is no assignment that is stable for $\widetilde{\Gamma}$ that assigns $i$ to $\mu(i)$. In other words, a student $i$ with information $\mathcal{H}_{i}$ has a legitimate complaint at an assignment $\mu$ if it is not possible to ‘rationalize’ $\mu(i)$.

### 29 Feb 2024 v2

@@ Abstract
In school districts where assignments are exclusively determined by a clearinghouse students can only appeal their assignment with a valid reason. An assignment is incontestable if it is appeal-proof. We study incontestability when students do not observe the other students’ preferences and assignments. Incontestability is shown to be equivalent to individual rationality, non-wastefulness, and respect for top-priority sets (a weakening of justified envy). Stable mechanisms and those Pareto dominating them are incontestable, as well as the Top-Trading Cycle mechanism (but Boston is not). Under a mild consistency property, incontestable mechanisms are $i$-indinstiguishable (Li,, 2017), and share similar incentive properties.
@@ 1 Introduction
Stability is a desirable objective when designing matching or assignment mechanisms because it prevents eventual disruptions by blocking pairs, a cause of market failure—see Roth, (1991) and Kagel and Roth, (2000). [^4] The rationale is easy to understand when considering a centralized job market, where workers can offer themselves to firms that rank higher in their preferences after being notified of their matches. If the matching were not stable, some firms might accept these ‘decentralized’ offers and thus reject candidates they have been assigned to by the clearinghouse. Incentives to participate in the matching mechanism are then weakened, which may end up in market failure. It should be noted, though, that stability does not always prevent market failure. See for instance McKinney et al., (2005)., [^5] Another common justification for stability is fairness. However, in school choice problems this normative criterion comes at the cost of efficiency —see (Ergin,, 2002). While this criterion is unquestionably needed in many cases, we contend that stability may be unnecessarily demanding in settings in which assignments are exclusively determined by a central authority, like in most school districts. In such markets, blocking pairs are not free to form independently, but arise through the actions of individual students who can appeal their assignment.
Appealing one’s assignment differ from the standard blocking concept because in practice most school districts only permit appeals by students who can provide justification for their complaint. For instance, a student could justify an appeal on the basis that given her priority ranking and school capacities, she should have gotten into, say, one of her top three schools, independently of the preferences of other students. In contrast, a student who is unhappy with her assignment but cannot provide a valid reason why her assignment should be different, has no recourse. What this means is that students may not have the information to identify every school with which they can form a blocking pair, so not all blocking pairs are a threat to an assignment. In other words, stability requires the absence of justified envy, but for a student to identify justified envy demands her to know the assignment of other students. This is usually not the case in school choice, especially in the first few weeks after notification of school assignment, when appeals must be made but students’ assignments being not (yet) public information. This is the point of departure of our paper: whether a school choice assignment is incontestable (appeal-proof) hinges on the information available to students.
The objective of this paper is to extend the stability concept by incorporating students’ ability to submit justified appeals. Specifically, we formalize the information students have access to when participating in a school choice mechanism, assuming that students only know their own assignment, and schools’ priority rankings, capacities and enrollment sizes. That is, students do not have information about the preferences and assignments of other students. We say that a student has a legitimate complaint if the information she has shows that her assignment cannot be the result of a stable assignment. In other words, a student has a legitimate complaint if her assignment is inconsistent with a stable assignment for any preference profile of the other students. An assignment with no legitimate complaint is called incontestable, and the paper studies such assignments.
@@
In principle, verifying incontestability of an assignment requires checking, for each student, all possible stable assignments given all possible preference profiles of the other students. Our first result (Theorem 1) shows that it can be verified much more easily: an assignment is incontestable if, and only if it is individually rational, non-wasteful, and satisfies a novel property, respect for top-priority sets. That latter requirement can be viewed as a non-justified envy condition adapted to the informational constraints of our setting. It differs from the standard non-justified envy in that it only applies to students for whom there exists a set of schools which they prefer to any other school and that cannot be filled only with students who have higher priority. We call such a set a top priority set, and an assignment respects top-priority sets if any student with such a set is assigned to a school in that set. When this property is violated for a student, one of the schools in her top-priority set must either be under capacity or have admitted a student with lower priority.
@@
Theorem 1 implies that a student has a legitimate complaint whenever: 1) she is not assigned to a school in her top priority set (if she has one), or 2) she is assigned to a school she finds unacceptable (i.e., violating individual rationality) or 3) a school preferred to her assignment is not full. Under a mild consistency property, called top-top consistency, we obtain a counterpart characterization of incontestable mechanism outcomes that parallels Theorem 1 (Theorem 2). Top-top consistency is a weakening of the standard consistency property that applies only for student-school pairs that are mutually ranked first.666 We show in the Appendix that the most common mechanisms considered in the literature are top-top consistent. Theorem 2 says that for students with a top-priority set, any school in that set can be the outcome of an incontestable and top-top consistent mechanism. That is, for any school in the set, and for any incontestable and top-top consistent mechanism, there exists a preference profile of other students under which she is assigned to that school by the mechanism. As for the students without a top-priority set, the set of possible outcomes is the set of all individually rational assignments. Remarkably, this characterization does not depend on the particular mechanism that is used. In other words, mechanisms that are incontestable and top-top consistent are all $i$-indistinguishable in the sense of Li, (2017).
Students’ rather limited information might make it reasonable to consider incontestable mechanisms rather than the more demanding stable mechanisms. A key question though, is what does this buy us? Crucially, incontestability is compatible with efficiency in a way that stability is not: while stable assignments may be Pareto dominated by unstable ones, any assignment that Pareto dominates an incontestable assignment is also incontestable. Since stable assignments are always incontestable, any mechanism that Pareto improves the student-optimal stable mechanism (SOSM) is also incontestable. Therefore, incontestable mechanisms can provide Pareto improvements on stable mechanisms. This has policy relevance, as our results provide an efficiency justification for using mechanisms that may not satisfy stability, but satisfy incontestability. For instance, the Boston mechanism does not satisfy incontestability (nor stability) but is used in school assignment due to its efficiency properties. Conversely, SOSM is stable but not efficient. Our result provides a rationale for incontestable mechanisms that can satisfy both efficiency and some weakened notion of stability, such as Kesten,’s (2010) Efficiency Adjusted Deferred Acceptance Mechanism (EADAM). Interestingly, there are also incontestable mechanisms that are not Pareto comparable to SOSM, like the Top Trading cycle mechanism and some of its variants.
@@
Our characterization in Theorem 2 turns out to be extremely useful to study (and compare) the incentive properties of incontestable mechanisms. Of course, such mechanisms are not guaranteed to be strategyproof. For instance, EADAM is not strategyproof. Reny, (2022) nevertheless finds that under EADAM, truth-telling is a maxmin optimal strategy. Theorem 2 makes it relatively easy to extend Reny’s theorem to any incontestable and top-top consistent mechanism. We also show that we can easily get results that compare the manipulability of mechanisms, when varying the maximal number of schools students are permitted to include in their submitted preference lists.
@@
Related literature   Our paper intersects with several branches of the matching and assignment literature. The first of these branches deals with proposals to obtain assignments that Pareto dominate the student-optimal stable mechanism. In broad strokes, those mechanisms allow for certain students to have justified envy. [^7] See Dur et al., (2019) who propose a family of mechanisms based on this principle. A key contribution in that direction is Kesten,’s (2010)’s EADAM algorithm. His mechanism consists of identifying ‘interrupters,’ that is, students who will make the student-optimal assignment inefficient because they prefer certain schools to their assignment (but that they eventually do not get under SOSM). In Kesten’s algorithm, students are first asked whether they want to waive their priorities at the schools for which they are interrupters. Ehlers and Morrill, (2020) and Reny, (2022) propose different rationales than Kesten’s, but end up with a mechanism that is outcome equivalent to EADAM (when all students waive their priorities). Our approach differs from these contributions in that our notion of blocking is entirely based on the information available to the students. It is the lack of complete information on students’ side that allows us to implement assignments that Pareto dominates the student-optimal assignment, and not a tolerance for some level of justified envy.
@@
We are not the first to propose a definition of stability that depends on the information available to participants. An early contribution in that respect is Chakraborty et al., (2010) who consider the problem where agents can infer the types of potential partners from the observation of the whole matching. Another contribution is Yenmez, (2013), who considers ex-ante and interim stability conditions. [^8] See also Ehlers and Massó, (2015). A notable contribution is Liu et al., (2014), who extend Crawford and Knoer,’s (1981) model where firms do not observe workers’ types. Our work is similar to Liu et al., in the sense that blocking and legitimate complaint are both based on a worst-case scenario principle. In their paper, a worker-firm pair blocks a matching if they can be both better off under any possible belief, and in our paper a legitimate complaint is a complaint that holds for any possible preference profile of the other students. However, our contribution is somehow orthogonal to Liu et al., (2014) or Chakraborty et al., (2010). They assume that firms observe the whole matching but not workers’ types, whereas students in our model observe schools’ priorities (i.e., their ‘types’) but not the complete assignment.
@@
Our paper also deals with the question of agent’s information when participating in a mechanism, and Li,’s (2017) contribution is obviously key to frame and interpret some of our results. Our results complement those of Möller, (2022), who analyzes conditions under which students can be guaranteed that the central authority has used the mechanism it initially announced. Möller,’s (2022) results and ours share some similarities but are obtained under different approaches. Möller,’s contribution is motivated by questioning the central authority’s credibility, while our initial motivation is to understand what stability entails in a low information environment. [^9] Möller, assumes that all students observe the complete assignments. See also for Hakimov and Raghavan, (2022) for another contribution in a similar vein.
Finally, Our paper also contributes to the literature that compares the manipulability of school choice mechanisms (Pathak and Sönmez,, 2013). Decerf and Van der Linden, (2021) compare the occurence of dominant strategies in various constrained mechanisms. Theorem Theorem 2 permits us to extend their comparisons results that involve (constrained) SOSM by identifying a class of mechanisms with the same occurence of dominant strategies as (constrained) SOSM.
@@ 3.1 Information and complaints
It will prove convenient to make the distinction between the information held at interim and ex-post stages of an assignment mechanism, that is, before or after the assignment mechanism has computed an assignment for all students. [^10] We will not consider in this paper the ex-ante stage (i.e., before students know their preferences over schools). We thus say that an information set $\mathcal{H}_{i}$ of a student $i$ is interim if $\succ_{i}\in\mathcal{H}_{i}$ and $\mathcal{H}_{i}$ does not contain any (partial) information about an assignment. The information set $\mathcal{H}_{i}$ is ex-post if $\succ_{i}\in\mathcal{H}_{i}$ and $\mathcal{H}_{i}$ contains some information about a realized assignment. [^11] Our approach to formalize students’ information is similar to that of Fernandez, (2018) or Chen and Möller, (2023). However, these two papers only consider ex-post information structures.
@@
A school choice problem $\Gamma=(I,S,\succ,r,q)$ is compatible with an interim information set $\mathcal{H}_{i}$ if there is no element in $\Gamma$ that contradicts an element in $\mathcal{H}_{i}$. A school choice problem $\Gamma=(I,S,\succ,r,q)$ and assignment $\mu$ are compatible with an ex-post information set $\mathcal{H}_{i}$ if there is no element in $\Gamma$ and $\mu$ that contradicts an element in $\mathcal{H}_{i}$.121212 We restrict in this paper to situations where each student is not informed about the other students’ preferences and assignments. So $\Gamma=(I,S,\succ^{\prime},r,q)$ is compatible with $\mathcal{H}_{i}=\{I,S,\succ_{i},r,q\}$ if $\succ_{i}^{\prime}=\succ_{i}$. Similarly, $\Gamma=(I,S,\succ^{\prime},r,q)$ and $\mu^{\prime}$ are compatible with $\mathcal{H}_{i}=\{I,S,\succ_{i},r,q,\mu(i),(|\mu(s)|)_{s\in S}\}$ if $\succ_{i}^{\prime}=\succ_{i}$, $\mu^{\prime}(i)=\mu(i)$, and $|\mu^{\prime}(s)|=|\mu(s)|$ for each $s\in S$.
@@
We say that a student $i$ with an ex-post information set $\mathcal{H}_{i}$ has a legitimate complaint at an assignment $\mu$ if $\mathcal{H}_{i}$ is enough to show that if $\mu$ were a stable assignment then $i$ should not be assigned to $\mu(i)$. To understand more formally what we mean by the fact that $\mathcal{H}_{i}$ is ‘enough’, consider its negation. If the information set $\mathcal{H}_{i}$ is ‘not enough’ for $i$’s complaint then it must be that whether $i$ can be assigned to $\mu(i)$ at a stable assignment depends on some elements not included in $\mathcal{H}_{i}$. That is, there are two problems, say, $\widehat{\Gamma}$ and $\widetilde{\Gamma}$, that are both compatible with $\mathcal{H}_{i}$, and there exists an assignment $\widehat{\mu}$ that is stable for $\widehat{\Gamma}$ such that $\widehat{\mu}(i)=\mu(i)$ but there is no assignment that is stable for $\widetilde{\Gamma}$ that assigns $i$ to $\mu(i)$. In other words, a student $i$ with information $\mathcal{H}_{i}$ has a legitimate complaint at an assignment $\mu$ if it is not possible to ‘rationalize’ $\mu(i)$.